Determine the potential at a point 0.50 m (a) form a +20 μC point charge, (b) form a -20 μC point charge.


1
Expert's answer
2022-03-30T03:32:06-0400
"V=k\frac{q}{r}"

(a)

"V=9*10^9*\frac{20*10^{-6}}{0.50}=360000\:\rm V"

(b)

"V=9*10^9*\frac{-20*10^{-6}}{0.50}=-360000\:\rm V"